OpenStudy (anonymous):

solve the system x-4y=1 -2x+5y=4

OpenStudy (zepp):

When it says 'solve the system', it means that you have to find the point where two functions cross each other. So the first step is to put y on one side, then you set these two functions equal to each other (since when x is equal for both of them, the y is the same as well)

OpenStudy (zepp):

Ok here: First step: put y on one side: y = stuffs for both equations

OpenStudy (australopithecus):

x-4y=1 -2x+5y=4 Solve for a variable im using x-4y=1 x = 1 + 4y sub that into the second equation as x -2(1 + 4y)+5y=4 solve for y sub y into first or second equation to solve for x

OpenStudy (australopithecus):

hope that helps

OpenStudy (zepp):

^